A ball is projected horizontally from the top of the hill with a velocity of 30ms-1 if it reaches the ground 5sec later, the height of the hill is?
1
Expert's answer
2019-07-11T09:31:35-0400

If a ball is projected horizontally, the motion time is dependent only on height.

The height of the hill

"H=\frac{gt^2}{2}=\frac{9.8\:\rm{m/s^2}\times (5\:\rm{s})^2}{2}=122.5\:\rm{m}"